OpenStudy (anonymous):

A high school graduating class is made up of 542 students. There are 156 more girls than boys. How many boys re there so is the problem set up b + (b+156) = 542 2b+156=542 2b+156-156= 542-156 2b= 386 2/2b= 386/2

OpenStudy (***[isuru]***):

yes!! the steps u r following r correct... except the last one it should be 2b/2 = 386/2 simplify it ... adnuwill get ur answer !!!

OpenStudy (jdoe0001):

ditto

OpenStudy (anonymous):

Okay, if anyone could answer why am I doing the b+ (b+156). I understand the "B+156" part but why am I doing the other b? Is it because b+156 gives me the total amount of the girls and the solo b is the amount of boys all together. Sorry, I can solve it it's just the whole why I am using this way.

OpenStudy (***[isuru]***):

yes! b+ 165 gives u the total amount of girls b = total amount of boys total amount of students = 365 = total amount of boys + total amount of girls 365 = b + (b + 365)